OpenStudy (mtalhahassan2):

can how we find the derivative of the following function 3sqrt x -1+x/x^4

OpenStudy (anonymous):

ill help but you also have to try.

OpenStudy (mtalhahassan2):

ok sure

OpenStudy (anonymous):

The answer is e3x√(32x−1/2) e3x(32x−1/2)

OpenStudy (anonymous):

Now I need a medal

OpenStudy (mtalhahassan2):

@Redneckk__Lifee but how you get that

OpenStudy (anonymous):

online duhh. and why did you ask my name you already know it.

OpenStudy (anonymous):

I'm pretty sure I'm waiting for a medal!!!!!!

OpenStudy (marigirl):

Okay so we are differentiating \[\sqrt[3]{(x-1)}+ \frac{ x }{ x^4 }\] Is that the equation? I want to make sure its correct before we begin

OpenStudy (marigirl):

@MTALHAHASSAN2

OpenStudy (mtalhahassan2):

@marigirl nope its wrong

OpenStudy (marigirl):

I meant- did i type your QUESTION correctly

OpenStudy (mtalhahassan2):



OpenStudy (firekat97):

so its \[f(x) = 3\sqrt{x} - \frac{ 1 + x }{ x^4 }\] ???

OpenStudy (mtalhahassan2):

nope

OpenStudy (firekat97):

no, I mean is that the question?

OpenStudy (firekat97):

@MTALHAHASSAN2

OpenStudy (mtalhahassan2):

yes it is

OpenStudy (mtalhahassan2):

ok wait a sec

OpenStudy (firekat97):

okay, so what you want to start off doing, is bringing the \[x^4\] to the top, can you show me how you may do this?

OpenStudy (mtalhahassan2):

@FireKat97 wait a sec

OpenStudy (firekat97):

okay just tag me back whenever you're ready :)

OpenStudy (mtalhahassan2):